The cost of producing a certain item consist of P102 per unit for labor and material cost and P54 per


unit for other variable cost. The fixed cost per month amounts to P850,000. The item is sold at P740 each.


a. Determine the break-even quantity per month.


b. How many units must be produced each month in order that the net profit equals the cost?


c. What is the net profit it for a production of 4,000 units per month, in pesos?

1
Expert's answer
2021-12-13T10:24:46-0500

Solution:

a.). Break even in units = fixed costs ÷ contribution margin


Contribution margin per unit = sales price per unit - variable costs

=740 - (102 + 54) = 740 - 156 = 584


BEP (units) = 850,000 ÷ 584=1455.48


BEP (units) = 1456 units


b.). Let x be units that must be produced each month in order that the net profit equals the cost

584x-850,000= 102x + 54x + 850,000

428x = 1,700,000

X = 3,972 units


c.). Net profit = Contribution margin - fixed costs

Contribution margin = 584 × 4000 = 2,336,000

Net profit = 2,336,000 - 850,000 = 1,486,000